Fermi National Laboratory
 

Operation Summary for November 16, 2009 - November 18, 2009

 

Monday Nov. 16

The day shift began with Tevatron (TeV) studies, with Operations monitoring the Recycler stash (122.3E10), with the Antiproton Source stacking (22.8E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments waiting for the next store.

During the day shift, Operations reported that a TeV cryo system power lead at sector A2 was found with too high a temperature. Mechanical technicians accessed the A-E sectors to investigate at 9 AM. This access lasted until mid-afternoon but did not fix the power lead problem. During the access many other TeV maintenance items were repaired.

During the evening shift, Operations reported that the second TeV access of the day for the power lead problem lasted from 4 PM to 8 PM, after which the TeV was turned on and left at high energy to verify the lead was working well enough to get by.

Tuesday Nov. 17

The midnight shift began with TeV studies, with Operations monitoring the stash (439.5E10), with Pbar stacking (54.1E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations put the TeV through a dry squeeze, orbit checks, and then into HEP shot setup. Operations established store 7364 at 4:30 AM with an average initial luminosity (AIL) of 268.91E30.

During the day shift, Operations terminated store 7364 at 3 PM and put the TeV into shot setup.

During the evening shift, Operations established store 7365 at 4:15 PM with an AIL of 271E30.

Wednesday Nov. 18

The midnight shift began with Operations monitoring store 7365 (111.01E30) and the stash (219.6E10), with Pbar stacking (21.0E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, at 2 AM, a Pbar transfer line power supply tripped off. This halted stacking. Since there would be an access in the morning, operators did not call in technicians to make repairs. Operations terminated store 7365 at 7 AM. Operators prepared the TeV for an access during the day shift. NuMI and MiniBooNE users took protons until 7 AM. NuMI reported that they had set a new record for the number of protons on target.

The Plans for Wednesday

The plans for today are for a six to eight hour access into the TeV to repair the sector A2 lead. MI and Recycler will remain up to maintain the stash. The Run Coordinator will allow other machines to access, perform maintenance, and conduct studies up to noon.
 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/18/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory